OpenStudy (anonymous):

you roll a standard number cube once. find P(0)

OpenStudy (jessiegonzales):

There are 4 numbers on a cube (3, 4, 5, 6) that are greater than 2. That's out of 6 possible rolls, so the probability is 4/6 = 2/3 For the coin, there is 1 head out of 2 outcomes = 1/2 Multiplying the two probabilities together you get: 2/3 x 1/2 = 1/3

OpenStudy (campbell_st):

well the problem depends on how many zeros are on your number cube as the the probability of P(0)

OpenStudy (anonymous):

i'm confused

OpenStudy (campbell_st):

well on your standard number cube... are there any zeros...?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

ok... so the probability of rolling a zero P(0) = 0 does that make sense...?

OpenStudy (anonymous):

yes